OpenStudy (anonymous):

Kevin sold 2 unicycles and 1 bicycle for a total of $110. Greg sold 4 unicycles and 3 bicycles for a total cost of $268. What is the cost of 1 bicycle? Answer $27 $31 $48 $79 Kevin sold 2 unicycles and 1 bicycle for a total of $110. Greg sold 4 unicycles and 3 bicycles for a total cost of $268. What is the cost of 1 bicycle? Answer $27 $31 $48 $79 @Mathematics

OpenStudy (anonymous):

2x + y = 110 4x + 3y = 268

OpenStudy (anonymous):

-3(2x + y = 110) 4x + 3y = 268

OpenStudy (anonymous):

-6x - 3y = -330 4x + 3y = 268

OpenStudy (anonymous):

-2x = -62 x = 31

OpenStudy (anonymous):

2(31) + y = 110 62 + y = 110 y = 48 $48 for 1 bicycle

OpenStudy (anonymous):

Ahhhhhhhhhh ty so much; this was corrcet :D